Hi, I am trying to solve this Linear Equation System by using the substitution method:
-x+3y=-1
x-2y=6

Answer by ichudov(499)

You can rewrite the second equation as saying
x=6+2y
then replace -(x) in the first equation with -(6+2y):
-(6+2y)+3y=-1
-6-2y+3y=-1
-6+y=-1
y=-1+6
y=5
Since x = 6+2y, x = 6+2*5 = 16.